# Rescue: Retrieval Augmented Secure Code Generation

### v1

@@ Abstract
Despite recent advances, Large Language Models (LLMs) still generate vulnerable code. Retrieval-Augmented Generation (RAG) has the potential to enhance LLMs for secure code generation by incorporating external security knowledge. However, the conventional RAG design struggles with the noise of raw security-related documents, and existing retrieval methods overlook the significant security semantics implicitly embedded in task descriptions. To address these issues, we propose Rescue, a new RAG framework for secure code generation with two key innovations. First, we propose a hybrid knowledge base construction method that combines LLM-assisted cluster-then-summarize distillation with program slicing, producing both high-level security guidelines and concise, security-focused code examples. Second, we design a hierarchical multi-faceted retrieval to traverse the constructed knowledge base from top to bottom and integrates multiple security-critical facts at each hierarchical level, ensuring comprehensive and accurate retrieval. We evaluated Rescue on four benchmarks and compared it with five state-of-the-art secure code generation methods on six LLMs. The results demonstrate that Rescue improves the SecurePass@1 metric by an average of 4.8 points, establishing a new state-of-the-art performance for security. Furthermore, we performed in-depth analysis and ablation studies to rigorously validate the effectiveness of individual components in Rescue.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities in coding-related tasks (Peng et al., 2023; Paradis et al., 2025). However, recent studies have revealed that LLMs often generate code with vulnerabilities  (Pearce et al., 2022; Fu et al., 2023; Majdinasab et al., 2024). He & Vechev (2023); He et al. (2024) propose to finetune LLMs with security-aware objects. Yet these methods require significant effort in data curation and finetuning. Constrained decoding mechanisms can prevent the model from generating insecure code without finetuning (Li et al., 2024a; Fu et al., 2024). However, they require the availability of trained security models or human-crafted rules to serve as oracles or constraints to detect insecure code tokens during decoding. Another line of research (Nazzal et al., 2024; Kim et al., 2024) leverages security analysis tools such as Bandit (PyCQA, 2025) and SpotBugs (SpotBugs, 2025) to provide vulnerability feedback for iterative code refinement. However, these security analysis tools heavily rely on pre-defined static analysis logic and heuristics to find bugs and vulnerabilities, which makes them less flexible to incorporate new vulnerabilities and security knowledge. Furthermore, since these security analysis tools are often used to assess the security of LLM-generated code in evaluation (Nazzal et al., 2024; Kim et al., 2024), using them as a verifier to provide feedback in the code generation stage raises a data leakage concern.
@@
Retrieval-Augmented Generation (RAG) offers a more flexible and training-free solution to incorporate security knowledge, such as documentation of secure coding practices and code examples. Despite some recent investigation (Zhang et al., 2024; Mukherjee & Hellendoorn, 2025), existing methods merely adapt conventional RAG methods to security domains and suffer from two limitations. First, security-related documents often contain information not relevant to the target coding task, which would unnecessarily distract the LLM from generating code for the target task. For instance, a code example that demonstrates a secure coding practice may contain code logic of an example task that is very different from the target task. Second, the retrievers used by existing methods simply treat all security-related information as general text and measure similarity between texts as security relevance. They do not capture security semantics, such as the security requirements implicitly embedded in a task description. This oversight leads to inaccurate retrieval of security-related data.
@@
To address these limitations, we propose Rescue, a REtrieval-augmented Secure Code gEneration framework. Rescue has two key innovations. First, we propose a hybrid knowledge base construction method that combines semantic summarization with static program analysis. Given raw security data, an LLM-assisted cluster-then-summarize pipeline first distills generalizable security knowledge as high-level guidelines (e.g.,“replace yaml.load() with yaml.safe_load() to prevent code execution vulnerabilities”). Then, a static program slicing procedure extracts concise, security-focused code examples by isolating the statements relevant to vulnerability fixes while filtering out unrelated logic. Second, we develop a hierarchical multi-faceted retrieval method that first performs a coarse-grained search to identify relevant vulnerability types and secure code examples. In each search iteration, our method proactively analyzes three security-critical aspects—API pattern, vulnerability cause analysis, and code—and then fuses these separated faceted search results to obtain precise security knowledge to guide code generation.
@@ 2.1 Hierarchical Security Knowledge Base Construction
Inspired by how human experts reason about security problems using taxonomies (Igure & Williams, 2008), we propose modeling security knowledge as a hierarchical structure, with generalizable knowledge at the high level and concise code examples at the low level. In particular, we leverage the Common Weakness Enumeration (CWE) (MITRE, 2025) to categorize vulnerabilities and construct this hierarchy.
In this work, we focus on constructing a security knowledge base from historical vulnerabilities and their fixes. Figure 2 shows an example. Such security data are widely available in many security databases (CVE Program, 2025; National Institute of Standards and Technology (2024), NIST; GitHub, 2024). Specifically, we use the training dataset from SafeCoder (He et al., 2024) as our raw security dataset in this work. This dataset includes 704 vulnerabilities and their fixes from CVE and GitHub projects. Please refer to Appendix A for more details.
@@ 2.1.1 Cluster-then-Summarize
Raw vulnerability-fix instances are often extensive and overly specific, making them unsuitable as direct general guidance. Although CWE has provided concise descriptions, they lack actionable fixing strategies and are too abstract to be effective for secure code generation. Prior work (He et al., 2024) demonstrates that using CWE descriptions in context does not significantly improve the security of LLM-generated code. To address this gap, we introduce a cluster-then-summarize pipeline to distill generalizable security knowledge from raw vulnerability-fix instances.
First, we group the raw data instances into clusters based on their associated CWE type. Then, for each cluster, we apply a bottom-up summarization process using an LLM. This process begins by summarizing small, fixed-size subsets of instances. Next, it recursively summarizes the outputs from the previous result until a single, cohesive summary is generated for the entire cluster. This hierarchical summarization approach allows us to effectively process a large volume of data while maintaining high-quality, comprehensive outputs. Details of the algorithm are provided in Appendix C.1.
Since this preprocessing step is a one-time effort, we adopt GPT-4o in an offline setting. In the end, this pipeline generates two summaries for each CWE category. First, it produces security guidelines that define actionable instructions and best practices for preventing specific vulnerabilities as high-level knowledge. Second, it summarizes vulnerability causes, which capture the root conditions and failure patterns that lead to the occurrence of the vulnerability. These distilled causes are subsequently leveraged in the retrieval stage as a key security aspect.
@@ 2.1.2 Security-Focused Static Program Slicing
Rescue begins by building a Program Dependence Graph (PDG) that captures how statements in the code depend on each other through both data and control dependence. From security patches, it identifies points of interest by treating deleted statements as indicators of vulnerable code and added statements as indicators of secure code. Around these points, we perform bidirectional slicing to extract the relevant context: backward slicing traces the statements that influence the vulnerable or secure code, while forward slicing captures those that are affected by it. Finally, the sliced subgraphs from the vulnerable and secure versions are compared and complemented with missing context, resulting in two parallel, contextually complete code variants. Appendix C.2 describes this security-focused slicing algorithm in detail.
@@ 2.2.1 Proactive Multi-Faceted Analysis
Conventional retrieval typically relies on task functional descriptions, which lack explicit security semantics. In contrast, our method proactively analyzes coding tasks from multiple security-critical perspectives. This proactive stance is essential because security vulnerabilities often emerge from subtle implementation details that are not apparent in task descriptions. Specifically, we consider three facets:
@@
(3) API Call Extraction: Since security vulnerabilities frequently stem from API misuse (Zhang et al., 2018; Li et al., 2021; Egele et al., 2013), we apply visitor pattern to traverse the abstract syntax trees of the draft code $C_{\text{draft}}$ to identify all API calls.
@@
Step 1: CWE-level Retrieval. Rescue first selects the top-k relevant CWE types using the first two facets: (1) Vulnerability Cause Analysis: Rescue uses a widely used dense retriever for RAG, bge-base-en-v1.5 (Xiao et al., 2023), to compute $\text{score}_{\text{VCA}}$ between the task’s vulnerability cause and the indexed vulnerability causes for each CWE type. (2) API Pattern: Since API calls are discrete, Rescue computes $\text{score}_{\text{API}}$ between the draft code’s APIs and all APIs associated with each CWE type via a sparse retrieval, BM25 (Robertson et al., 1995).
@@
Step 2: Code-level Retrieval. After narrowing down to relevant CWE types, we proceed to conduct a fine-grained search at the lower level using the same two facets and an additional third facet, Code Similarity. We also employ dense retrieval to obtain $\text{score}_{\text{C}}$ between the draft code and sliced secure code examples, capturing similarity at the code level to identify relevant secure patterns.
@@ 3.1 Experiment Setup
Benchmarks We first evaluate Rescue on a state-of-the-art security benchmark called CodeGuard+ (Fu et al., 2024). CodeGuard+ incorporates and extends prior security benchmarks (Pearce et al., 2022; Siddiq & Santos, 2022). It includes 94 security-sensitive coding scenarios, each of which is accompanied by unit tests and CodeQL checks, enabling the testing of both functional correctness and security. CodeQL (GitHub, 2025) is a widely used security analysis tool and has been used to evaluate the security of LLM-generated code in several studies (He & Vechev, 2023; He et al., 2024; Li et al., 2024a).
@@
Metrics Following prior work (Zhang et al., 2024; He et al., 2024; Li et al., 2024a; He & Vechev, 2023), we used SecureRate to measure the security of LLM-generated code. SecureRate is defined as the proportion of generated code samples that pass the security checks, e.g., CodeQL checks in CodeGuard+ (Fu et al., 2024). For functional correctness measurement, we followed recent work (Zhuo et al., 2025; Jain et al., 2025) to use the unbiased $\text{Pass}@k$ (Chen et al., 2021).
@@
Baseline Methods We compare our approach with five state-of-the-art baseline methods, each representing a different strategy for secure code generation: (1) SecCoder (Zhang et al., 2024) is a RAG method that retrieves the most similar secure code example to facilitate in-context learning. To ensure a fair comparison, we utilize the same retriever bge-base-en-v1.5 (Xiao et al., 2023) and the raw security data introduced in Section 2 as retrieval documents. (2) Codexity (Kim et al., 2024) utilizes an external security tool to identify vulnerabilities for generated code and then iteratively refines the code based on detection results, up to three iterations. To avoid data leakage, we use another widely adopted security tool, semgrep (Semgrep, 2025), instead of CodeQL. (3) SafeCoder (He et al., 2024) adopts a specialized instruction-tuning method for code security. We use the training dataset and hyperparameters from the original paper to finetune the LLMs used in our experiments. Given the limited computational resources, we only fine-tune Qwen2.5-Coder-7B-Instruct, Llama3.1-8B-Instruct, and Deepseek-Coder-V2-Lite-Instruct, excluding larger models and closed-source models. (4) CoSec (Li et al., 2024a) manipulates the decoding process by controlling the output token logits based on a small trained security-specialized model. For implementation, we use the same training dataset as in the original paper. Since this method requires training a smaller model that uses the same tokenizer to perform its decoding, we select Qwen2.5-Coder-0.5B-Instruct as the small security-specialized model for Qwen2.5-Coder-7B and 32B and Llama3.1-1B-Instruct for Llama3.1-8B. Other models are either closed-source or lack a smaller version from the same model family. (5) INDICT (Le et al., 2024) is a multi-agent debate framework with external security tools to generate both security and helpfulness critiques for iterative code refinement. Following the original setup, we set the iteration round to three and use all four tools for code generation.
Other Implementation Details We use tree-sitter (Brunsfeld & Github, ) to implement static program analysis, slicing, and API extraction. To control the length of sliced code, our method performs 2-hop program slicing. During generation, we set the temperature to 0.2. To balance the precision and recall, we use a top-k value of 4. To only search security knowledge for security-relevant problems, the thresholds for the facets of API, vulnerability cause, and code are set to 4.0, 0.75, and 0.65, respectively. We follow prior work (Cormack et al., 2009) to set the RRF parameter $\alpha$ to 60.
@@ 3.2 Main Results
As shown in Table 1, Rescue consistently outperforms all existing methods in terms of SecurePass@1, the metric that balances security and functional correctness of generated code. Specifically, Rescue achieves 4.8% absolute improvement on average compared with the second-best baseline. We also observed that several existing approaches sacrifice functional correctness for security. For example, though INDICT (Le et al., 2024) achieves a higher SecureRate than Rescue, it has the lowest Pass@1 among all methods. Furthermore, when evaluated on the three benchmarks on functional correctness, Rescue achieves comparable performance compared with the original models. This indicates that Rescue does not severely damage the functional correctness of generated code while improving its security. Appendix B.1 provides a more direct comparison of the performance improvements achieved by each method.

### v2

@@ Abstract
Despite recent advances, Large Language Models (LLMs) still generate vulnerable code. Retrieval-Augmented Generation (RAG) has the potential to enhance LLMs for secure code generation by incorporating external security knowledge. However, the conventional RAG design struggles with the noise of raw security-related documents, and existing retrieval methods overlook the significant security semantics implicitly embedded in task descriptions. To address these issues, we propose Rescue, a new RAG framework for secure code generation with two key innovations. First, we propose a hybrid knowledge base construction method that combines LLM-assisted cluster-then-summarize distillation with program slicing, producing both high-level security guidelines and concise, security-focused code examples. Second, we design a hierarchical multi-faceted retrieval that traverses the constructed knowledge base from top to bottom and integrates multiple security-critical facts at each hierarchical level, ensuring comprehensive and accurate retrieval. We evaluated Rescue on four benchmarks and compared it with five state-of-the-art secure code generation methods on six LLMs. The results demonstrate that Rescue improves the SecurePass@1 metric by an average of 4.8 points, establishing a new state-of-the-art performance for security. Furthermore, we performed in-depth analysis and ablation studies to rigorously validate the effectiveness of individual components in Rescue. Our code is available at https://github.com/steven1518/RESCUE.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities in coding-related tasks (Peng et al., 2023; Paradis et al., 2025). However, recent studies have revealed that LLMs often generate code with vulnerabilities  (Pearce et al., 2022; Fu et al., 2023; Majdinasab et al., 2024). He and Vechev (2023); He et al. (2024) propose to finetune LLMs with security-aware objects. Yet these methods require significant effort in data curation and finetuning. Constrained decoding mechanisms can prevent the model from generating insecure code without finetuning (Li et al., 2024a; Fu et al., 2024). However, they require the availability of trained security models or human-crafted rules to serve as oracles or constraints to detect insecure code tokens during decoding. Another line of research (Nazzal et al., 2024; Kim et al., 2024) leverages security analysis tools such as Bandit (PyCQA, 2025) and SpotBugs (SpotBugs, 2025) to provide vulnerability feedback for iterative code refinement. However, these security analysis tools heavily rely on pre-defined static analysis logic and heuristics to find bugs and vulnerabilities, which makes them less flexible to incorporate new vulnerabilities and security knowledge. Furthermore, since these security analysis tools are often used to assess the security of LLM-generated code in evaluation (Nazzal et al., 2024; Kim et al., 2024), using them as a verifier to provide feedback in the code generation stage raises a data leakage concern.
@@
Retrieval-Augmented Generation (RAG) offers a more flexible and training-free solution to incorporate security knowledge, such as documentation of secure coding practices and code examples. Despite some recent investigation (Zhang et al., 2024; Mukherjee and Hellendoorn, 2025; Tony et al., 2025), existing methods merely adapt conventional RAG methods to security domains and suffer from two limitations. First, security-related documents often contain information not relevant to the target coding task, which would unnecessarily distract the LLM from generating code for the target task. For instance, a code example that demonstrates a secure coding practice may contain code logic of an example task that is very different from the target task. Second, the retrievers used by existing methods simply treat all security-related information as general text and measure similarity between texts as security relevance. They do not capture security semantics, such as the security requirements implicitly embedded in a task description. This oversight leads to inaccurate retrieval of security-related data.
@@
To address these limitations, we propose Rescue, a REtrieval-augmented Secure Code gEneration framework. Rescue has two key innovations. First, we propose a hybrid knowledge base construction method that combines semantic summarization with static program analysis. Given raw security data, an LLM-assisted cluster-then-summarize pipeline first distills generalizable security knowledge as high-level guidelines (e.g.,“replace yaml.load() with yaml.safe_load() to prevent code execution vulnerabilities”). Then, a static program slicing procedure extracts concise, security-focused code examples by isolating the statements relevant to vulnerability fixes while filtering out unrelated logic. Second, we develop a hierarchical multi-faceted retrieval method that performs a coarse-grained search to identify relevant vulnerability types and secure code examples. In each search iteration, our method proactively analyzes three security-critical aspects—API pattern, vulnerability cause analysis, and code—and then fuses these separated faceted search results to obtain precise security knowledge to guide code generation.
@@ 2.1 Hierarchical Security Knowledge Base Construction
Inspired by how human experts reason about security problems using taxonomies (Igure and Williams, 2008), we propose modeling security knowledge as a hierarchical structure, with generalizable knowledge at the high level and concise code examples at the low level. In particular, we leverage the Common Weakness Enumeration (CWE) (MITRE, 2025) to categorize vulnerabilities and construct this hierarchy.
In this work, we focus on constructing a security knowledge base from historical vulnerabilities and their fixes. Figure 2 shows an example. Such security data are widely available in many security databases (CVE Program, 2025; National Institute of Standards and Technology (NIST), 2024; GitHub, 2024b). Specifically, we use the training dataset from SafeCoder (He et al., 2024) as our raw security dataset in this work. This dataset includes 704 vulnerabilities and their fixes from CVE and GitHub projects. Please refer to Appendix A for more details.
@@ 2.1.1 Cluster-then-Summarize
Raw vulnerability-fix instances are often verbose and instance-specific, making them unsuitable as direct general guidance. Although CWE has provided concise descriptions, they lack actionable fixing strategies and are too abstract to be effective for secure code generation. Prior work (He et al., 2024) demonstrates that using CWE descriptions in context does not significantly improve the security of LLM-generated code. To address this gap, we introduce a cluster-then-summarize pipeline to distill generalizable security knowledge from raw vulnerability-fix instances.
First, we group the raw data instances into clusters based on their associated CWE type. Then, for each cluster, we apply a bottom-up summarization process using an LLM. This process begins by summarizing small, fixed-size subsets of instances. Next, it recursively summarizes the outputs from the previous results until a single, cohesive summary is generated for the entire cluster. This recursive, bottom-up summarization approach allows us to effectively process a large volume of data while maintaining high-quality, comprehensive outputs. Details of the algorithm are provided in Appendix B.1.
Since this preprocessing step is a one-time effort, we adopt GPT-4o in an offline setting. Ultimately, this pipeline generates two summaries for each CWE category. First, it produces security guidelines that define actionable instructions and best practices for preventing specific vulnerabilities as high-level knowledge. Second, it summarizes vulnerability causes, which capture the root conditions and failure patterns that lead to the occurrence of the vulnerability. These distilled causes are subsequently leveraged in the retrieval stage as a key security aspect.
@@ 2.1.2 Security-Focused Static Program Slicing
Rescue begins by building a Program Dependence Graph (PDG) that captures how statements in the code depend on each other through both data and control dependence. Based on the statements influenced by patches, it identifies points of interest by treating deleted statements as indicators of vulnerable code and added statements as indicators of secure code. Around these points, we perform bidirectional slicing to extract the relevant context: backward slicing traces the statements that influence the vulnerable or secure code, while forward slicing captures those that are affected by it. Finally, the sliced subgraphs from the vulnerable and secure versions are compared and complemented with missing context, resulting in two parallel, contextually complete code variants. Appendix B.2 describes this security-focused slicing algorithm in detail.
@@ 2.2.1 Proactive Multi-Faceted Analysis
Conventional retrieval typically relies on task functional descriptions, which lack explicit security semantics. In contrast, our method proactively analyzes coding tasks from multiple security-critical perspectives. This proactive stance is essential because security vulnerabilities often emerge from subtle implementation details that are not apparent in task descriptions. Specifically, we consider the following three facets:
@@
(3) API Call Extraction: Since security vulnerabilities frequently stem from API misuse (Zhang et al., 2018; Li et al., 2021; Egele et al., 2013), we apply the visitor pattern to traverse the abstract syntax trees of the draft code $C_{\text{draft}}$ to identify all API calls.
@@
Step 1: CWE-level Retrieval. Rescue first selects the top-k relevant CWE types using two facets: (1) Vulnerability Cause Analysis: Rescue uses a widely used dense retriever for RAG, bge-base-en-v1.5 (Xiao et al., 2023), to compute $\text{score}_{\text{VCA}}$ between input task’s vulnerability cause and the indexed vulnerability causes per CWE type. These per-CWE vulnerability causes are distilled from raw instances during the offline stage (Section 2.1.1). (2) API Pattern: Since API calls are discrete, Rescue computes $\text{score}_{\text{API}}$ between the draft code’s APIs and the API set of each CWE type via sparse retrieval, BM25 (Robertson et al., 1995). Each CWE’s API set is constructed by aggregating all APIs extracted from its associated vulnerability-fix instances.
@@
Step 2: Code-level Retrieval. After narrowing down to relevant CWE types, we proceed to conduct a fine-grained search at the low level using the same two facets and an additional third facet, Code Similarity. We also employ dense retrieval to obtain $\text{score}_{\text{C}}$ between the draft code and sliced secure code examples, capturing similarity at the code level to identify relevant secure patterns.
@@ 3.1 Experiment Setup
Benchmarks We first evaluate Rescue on a state-of-the-art security benchmark called CodeGuard+ (Fu et al., 2024). CodeGuard+ incorporates and extends prior security benchmarks (Pearce et al., 2022; Siddiq and Santos, 2022). It includes 94 security-sensitive coding scenarios, each of which is accompanied by unit tests and CodeQL checks, enabling the testing of both functional correctness and security. CodeQL (GitHub, 2025) is a widely used security analysis tool and has been used to evaluate the security of LLM-generated code in several studies (He and Vechev, 2023; He et al., 2024; Li et al., 2024a).
@@
Metrics Following prior work (Zhang et al., 2024; He et al., 2024; Li et al., 2024a; He and Vechev, 2023), we used SecureRate to measure the security of LLM-generated code. SecureRate is defined as the proportion of generated code samples that pass the security checks, e.g., CodeQL checks in CodeGuard+ (Fu et al., 2024). For functional correctness measurement, we followed recent work (Zhuo et al., 2025; Jain et al., 2025) to use the unbiased $\text{Pass}@k$ (Chen et al., 2021).
@@
Baseline Methods We compare our approach with five state-of-the-art baseline methods, each representing a different strategy for secure code generation: (1) SecCoder (Zhang et al., 2024) is a RAG method that retrieves the most similar secure code example to facilitate in-context learning. To ensure a fair comparison, we utilize the same retriever bge-base-en-v1.5 (Xiao et al., 2023) and the raw security data introduced in Section 2 as retrieval documents. (2) Codexity (Kim et al., 2024) utilizes an external security tool to identify vulnerabilities for generated code and then iteratively refines the code based on detection results, up to three iterations. To avoid data leakage, we use another widely adopted security tool, semgrep (Semgrep, 2025), instead of CodeQL. (3) SafeCoder (He et al., 2024) adopts a specialized instruction-tuning method for code security. We use the training dataset and hyperparameters from the original paper to finetune the LLMs used in our experiments. Given the limited computational resources, we only fine-tune Qwen2.5-Coder-7B-Instruct, Llama3.1-8B-Instruct, and Deepseek-Coder-V2-Lite-Instruct, excluding larger models and closed-source models. (4) CoSec (Li et al., 2024a) manipulates the decoding process by controlling the output token logits based on a small trained security-specialized model. For implementation, we use the same training dataset as in the original paper. Since this method requires training a smaller model that uses the same tokenizer to perform its decoding, we select Qwen2.5-Coder-0.5B-Instruct as the small security-specialized model for Qwen2.5-Coder-7B-Instruct and 32B and Llama3.1-1B-Instruct for Llama3.1-8B-Instruct. Other models are either closed-source or lack a smaller version from the same model family. (5) INDICT (Le et al., 2024) is a multi-agent debate framework with external security tools to generate both security and helpfulness critiques for iterative code refinement. Following the original setup, we set the iteration round to three and use all four tools for code generation.
Other Implementation Details We use tree-sitter (Brunsfeld and Github, ) to implement static program analysis, slicing, and API extraction. To control the length of sliced code, our method performs 2-hop program slicing. During generation, we set the temperature to 0.2. To balance the precision and recall, we use a top-k value of 4. To only search security knowledge for security-relevant problems, the thresholds for the facets of API, vulnerability cause, and code are set to 4.0, 0.75, and 0.65, respectively. We follow prior work (Cormack et al., 2009) to set the RRF parameter $\alpha$ to 60.
@@ 3.2 Main Results
As shown in Table 1, Rescue consistently outperforms all existing methods in terms of SecurePass@1, the metric that balances security and functional correctness of generated code. Specifically, Rescue achieves 4.8% absolute improvement on average compared with the second-best baseline. We also observed that several existing approaches sacrifice functional correctness for security. For example, though INDICT (Le et al., 2024) achieves a higher SecureRate than Rescue, it has the lowest Pass@1 among all methods. Furthermore, when evaluated on the three benchmarks on functional correctness, Rescue achieves comparable performance compared with the original models. This indicates that Rescue does not severely damage the functional correctness of generated code while improving its security. Appendix C.1 provides a more direct comparison of the performance improvements achieved by each method.